You are given the following data.
Number of Absences Final Grade

0 98

1 91

2 76

3 61

4 60

5 58


- Find the correlation coefficient for the data.
- Find the equation for the regression line for the data, and predict the final grade of a student who misses 3.5 days.
For the number of absenses and final grade, they are side by side.
Example:
absence 0 = grade of 98
absence 1 = grade of 91.
absence 2 = grade of 76, etc

Answer by ewatrrr(24785)

 
Hi,
My recommendation would be that You become familiar with either a TI Calculator
or Excel functions/Stat Tool for making Your statistical calculations.
For ex:
using Excel CORREL function:
entering Data Given...
correlation coefficient = -.9537
Using www.easycalculation.com/statistics/regression.php
Entering Data Given...
y = 96-8.8x
final grade of a student who misses 3.5 days
y = 96-8.8(3.5) = 65.2